OpenStudy (anonymous):

how do i reverse the signs when verifying that tan (x+pi/2)= -cotx

OpenStudy (anonymous):

My approach could be wrong, but Im thinking this: \[\tan(x+\frac{ \pi }{ 2 }) = \tan[-(-x-\frac{ \pi }{ 2 })]\] Because tangent is odd function, we can say: \[\tan(-\theta) = -\tan \theta\]That takes care of the negative out front. Also, the period of tangent is pi, so -pi/2 would have the same value for tangent is pi/2. So rewriting -pi/2 as pi/2 is no problem: \[-\tan(-x-\frac{ \pi }{ 2 }) \implies -\tan(\frac{ \pi }{ 2 }-x) = -cotx\]

OpenStudy (anonymous):

thank you!!!!!

OpenStudy (anonymous):

Np :3

OpenStudy (anonymous):

Do you know All,silver,tea,cups formula...

OpenStudy (anonymous):

no

OpenStudy (anonymous):



OpenStudy (phi):

If I get stuck, I go back to basics. tan(x) = sin(x)/cos(x) tan(a+b) = sin(a+b)/cos(a+b) with b=90º and the sum of angles formula sin(a+90)= sin(a) cos(90) + cos(a) sin(90) which simplifies to cos(a) cos(a+90)= cos(a) cos(90) - sin(a) sin(90) which simplifies to -sin(a) and cos(a)/-sin(a) = -cot(a)

OpenStudy (anonymous):

Haha, I like that one.

OpenStudy (anonymous):

Tan(pi/2+x) means it is in 2nd quadrent, so it becomes -ve

OpenStudy (anonymous):

ohhhh